# Spatial disaggregation of time series

### v1

@@ Abstract
Spatial models , Spatial autoregressive model , heteroscedasticity , Spatial regression , Maximun likelihood estimation
@@ 1 Introduction
Spatiotemporal data modeling has gained increasing relevance in econometrics and regional science due to the simultaneous presence of spatial spillovers and temporal dynamics in key macroeconomic indicators. Spatial autoregressive (SAR) models have long been used to capture spatial autocorrelation between neighboring regions, improving inference and prediction by explicitly considering spatial dependencies (LeSage and Pace, 2015; Cressie and Wikle, 2011). Recent developments have extended these models to incorporate temporal structure, leading to comprehensive spatiotemporal frameworks that model both forms of dependence jointly (Gelfand et al., 2019; Hsu et al., 2021).
@@
This article contributes to the literature by proposing a new methodology for the spatiotemporal disaggregation of economic aggregates, combining spatial autoregressive models, benchmarking restrictions, and covariate information. Before presenting our approach, it is important to highlight several features that make it attractive both theoretically and practically. First, the method is based on a rigorous asymptotic theory: we establish global identifiability and asymptotic normality of the joint estimator under general conditions that do not require Gaussianity. This makes the model robust to the nonnormal and heteroskedastic nature of economic residuals. Second, we introduce a novel anchoring mechanism that incorporates partially observed regional values as linear restrictions. This step mimics real-life situations with sporadic regional data (e.g., for capital cities or large communities) and helps reduce the variance of the predictions. Third, we conducted an exhaustive simulation study encompassing 74,088 parameter configurations, covering a wide range of spatial autocorrelations, covariate informativeness, and grid sizes. These simulations confirmed the robustness and stability of our approach across various scenarios. Finally, we demonstrated the practical relevance of our method with a case study on the disaggregation of Spanish GDP into the 17 autonomous communities between 2002 and 2023, using a large set of auxiliary indicators and principal component analysis for dimensionality reduction.
@@ 2 Spatial autoregressive model
Let $Y_{t}$ be a time series measured at time $t=1,\ldots,T$ over the spatial unit $\mathcal{R}$, which is subdivided in $n$ spatial regions $\{\mathcal{R}_{1},\ldots,\mathcal{R}_{n}\}$; let $\mathbf{z}_{it}$ a set of $k$ known explicative variables measured at time $t$ in spatial region $i$. Let $Y_{it}$ the $i$th unknown time serie disagregated at time $t$ over the region $i$ with the restriction defined as:
$\sum_{i=1}^{n}Y_{it}=Y_{t},\qquad\forall t=1,\ldots,T$ (1)
@@
The objective is obtain a estimated values for $Y_{it}$ using the information allocate both in $\mathbf{z}_{it}$ and $\mathbf{Y}_{a}=(Y_{1},\ldots,Y_{T})$. A spatial autoregressive model is supposed for the vector $\mathbf{Y}_{t}=(Y_{1t},\ldots,Y_{nt})$, the completed model is given by:
$Y_{it}=\rho\sum_{j=1}^{n}w_{ij}Y_{jt}+\mathbf{z}_{it}^{\top}\boldsymbol{\beta}+u_{it}$ (2)
where $\epsilon_{t}\sim N(0,\sigma^{2})$, $|\rho|<1$, and
$\mathbf{Z}$ $=\begin{bmatrix}\mathbf{z}_{11}^{\top}\\ \mathbf{z}_{21}^{\top}\\ \vdots\\ \mathbf{z}_{n1}^{\top}\\ \mathbf{z}_{12}^{\top}\\ \vdots\\ \mathbf{z}_{n2}^{\top}\\ \vdots\\ \mathbf{z}_{nT}^{\top}\\ \end{bmatrix}_{nT\times k},\;\mathbf{Y}=\begin{bmatrix}\mathbf{Y}_{11}\\ \mathbf{Y}_{21}\\ \vdots\\ \mathbf{Y}_{n1}\\ \mathbf{Y}_{12}\\ \vdots\\ \mathbf{Y}_{n2}\\ \vdots\\ \mathbf{Y}_{nT}\\ \end{bmatrix}_{nT\times 1},\;\mathbf{W}=\begin{bmatrix}w_{11}&w_{12}&\cdots&w_{1n}\\ w_{21}&w_{22}&\cdots&w_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ w_{n1}&w_{n2}&\cdots&w_{nn}\\ \end{bmatrix},\;\mathbf{U}=\begin{bmatrix}\mathbf{u}_{11}\\ \mathbf{u}_{21}\\ \vdots\\ \mathbf{u}_{n1}\\ \mathbf{u}_{12}\\ \vdots\\ \mathbf{u}_{n2}\\ \vdots\\ \mathbf{u}_{nT}\\ \end{bmatrix}_{nT\times 1}$ (3)
@@
For the vector $\mathbf{U}_{i}=(u_{i1},\ldots,u_{iT})$, an $ARIMA(1,0,0)$, i.e. $\phi_{1}(B)u_{it}=\epsilon_{it}$, and $\epsilon_{it}\sim(0,\sigma^{2})$ is proposed for modelling the unobservable time series, similar to the proposal of proietti2006temporal, Quilis (2018) and Bisio (2024). This assumption over $\mathbf{U}_{i}$ allow defined the covariance matrix of $\mathbf{U}_{i}$ as:
$\boldsymbol{\Sigma}_{U}=Var(\mathbf{U}_{i})=\left\{\frac{\sigma^{2}\phi_{1}^{|i-j|}}{1-\phi_{1}^{2}}\right\}_{T\times T}$ (4)
where $\phi$ is the autoregressive parameter of the $ARIMA(1,0,0)$ and $\sigma^{2}$ their associated variance. Then, the model defined in Equation (2) can be expressed as:
$\mathbf{Y}$ $=\rho(\mathbb{I}_{T}\otimes\mathbf{W})\mathbf{Y}+\mathbf{Z}\boldsymbol{\beta}+\mathbf{U}$ (5)
@@
the problem is that $\mathbf{Y}$ and $\mathbf{U}$ are unobservable by the characteristics of the measure. But $\mathbf{Z}$ and $Y_{t}$ are observable and known, then, defined the matrix $\mathbf{C}$ as:
$\mathbf{C}=\mathbb{I}_{T}\otimes\mathbf{1}_{n}^{\top}$ (6)
such that $\mathbf{CY}=\mathbf{Y}_{a}$, then Equation (5) can be expressed as:
$\mathbf{Y}$ $=[\mathbf{I}_{nT}-\rho(\mathbf{I}_{T}\otimes\mathbf{W})]^{-1}\mathbf{Z}\boldsymbol{\beta}+[\mathbf{I}_{nT}-\rho(\mathbf{I}_{T}\otimes\mathbf{W})]^{-1}\mathbf{U}$
$\mathbf{CY}$ $=\mathbf{C}[\mathbf{I}_{nT}-\rho(\mathbf{I}_{T}\otimes\mathbf{W})]^{-1}\mathbf{Z}\boldsymbol{\beta}+\mathbf{C}[\mathbf{I}_{nT}-\rho(\mathbf{I}_{T}\otimes\mathbf{W})]^{-1}\mathbf{U}$
$\mathbf{Y}_{a}$ $=\mathbf{C}[\mathbf{I}_{T}\otimes(\mathbf{I}_{n}-\rho\mathbf{W})]^{-1}\mathbf{Z}\boldsymbol{\beta}+\mathbf{C}[\mathbf{I}_{T}\otimes(\mathbf{I}_{n}-\rho\mathbf{W})]^{-1}\mathbf{U}$
$\mathbf{Y}_{a}$ $=\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}+\mathbf{C}\mathbf{A}^{-1}\mathbf{U}$ (7)
It is obtained that:
$Cov(\mathbf{U})_{nT\times nT}$ $=\boldsymbol{\Sigma}_{U}\otimes\mathbb{I}_{n}$
$Cov(\mathbf{C}\mathbf{A}^{-1}\mathbf{U})_{T\times T}$ $=\mathbf{C}\mathbf{A}^{-1}(\boldsymbol{\Sigma}_{U}\otimes\mathbb{I}_{n})(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}=\boldsymbol{\Sigma}_{Y}=\mathbf{B}$ (8)
therefore, the generalized least squares estimator for $\boldsymbol{\beta}$ is given by:
$\hat{\boldsymbol{\beta}}$ $=(\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\mathbf{B}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z})^{-1}\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\mathbf{B}^{-1}\mathbf{Y}_{a}$ (9)
It should be noted that the estimator $\hat{\boldsymbol{\beta}}$ depends on the values of the parameters $\rho$, $\sigma^{2}$, and $\phi_{1}$. To estimate these parameters, the following quasi-likelihood maximization problem is solved:
$(\hat{\rho},\hat{\sigma}^{2},\hat{\phi}_{1})=\underset{(\rho,\phi_{1},\sigma^{2})\in\mathcal{U}}{\operatorname{argmax}}\Bigg{\{}-\frac{1}{2}\log|\boldsymbol{\Sigma}_{Y}|-\frac{1}{2}(\mathbf{Y}_{a}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}_{Y}^{-1}(\mathbf{Y}_{a}-\boldsymbol{\mu})\Bigg{\}}$ (10)
where
$\boldsymbol{\mu}=\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}\quad\text{and}\quad\boldsymbol{\Sigma}_{Y}=\mathbf{C}\mathbf{A}^{-1}(\boldsymbol{\Sigma}_{U}\otimes\mathbb{I}_{n})(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}.$
In practice, this maximization in Equation (10) is carried out in R Core Team (2024) using the optim function with the ”L-BFGS-B” algorithm, which allows the constraints on the parameter space $\mathcal{U}$ to be explicitly imposed. The L-BFGS-B algorithm is a limited-memory version of the Broyden–Fletcher–Goldfarb–Shanno (BFGS) quasi-Newton method, adapted for large-scale problems with simple box constraints (Byrd et al., 1995).
Based on this joint estimation approach, the following iterative procedure is proposed to obtain consistent estimators of $(\boldsymbol{\beta},\rho,\sigma^{2},\phi_{1})$ and the disaggregated spatio-temporal series:
  1. 1.
    Estimate $(\hat{\rho},\hat{\phi}_{1},\hat{\sigma}^{2})$ by maximizing the quasi-likelihood function using the optim function in R with the "L-BFGS-B" algorithm, which allows the box constraints $(\rho,\phi_{1})\in(-1,1)$ and $\sigma^{2}>0$ to be imposed explicitly (Byrd et al., 1995).
  2. 2.
    Given these estimates, obtain $\hat{\boldsymbol{\beta}}$ as:
    $\hat{\boldsymbol{\beta}}=\Big{(}\mathbf{Z}^{\top}\mathbf{A}^{-1}\mathbf{C}^{\top}(\mathbf{C}\mathbf{B}\mathbf{C}^{\top})^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\Big{)}^{-1}\mathbf{Z}^{\top}\mathbf{A}^{-1}\mathbf{C}^{\top}(\mathbf{C}\mathbf{B}\mathbf{C}^{\top})^{-1}\mathbf{Y}_{a}.$
  3. 3.
    The disaggregated series is then estimated by:
    $\widehat{\boldsymbol{Y}}=\widehat{\mathbf{A}}^{-1}\mathbf{Z}\hat{\boldsymbol{\beta}}+\widehat{\mathbf{B}}\,\mathbf{C}^{\top}\big{(}\mathbf{C}\widehat{\mathbf{B}}\,\mathbf{C}^{\top}\big{)}^{-1}\Big{(}\mathbf{Y}_{a}-\mathbf{C}\widehat{\mathbf{A}}^{-1}\mathbf{Z}\hat{\boldsymbol{\beta}}\Big{)}$.
  4. 4.
    Anchorage step. When partial information is available for specific regions and time points, such information can be incorporated as additional linear constraints of the form
    $\mathbf{H}\boldsymbol{Y}=\mathbf{d},$
    where $\mathbf{H}$ is a known indicator matrix (with entries in $\{0,1\}$) that selects the positions in $\boldsymbol{Y}$ for which values are known, and $\mathbf{d}$ is the corresponding vector of observed values. This anchorage does not require full annual coverage and may include only a few scattered observations across time and regions (e.g., benchmark values or administrative figures). The constraint is incorporated jointly with the aggregation constraint $\mathbf{C}\boldsymbol{Y}=\mathbf{Y}_{a}$ by stacking both into an extended constraint system:
    $\begin{bmatrix}\mathbf{C}\\ \mathbf{H}\end{bmatrix}\boldsymbol{Y}=\begin{bmatrix}\mathbf{Y}_{a}\\ \mathbf{d}\end{bmatrix}$.
    The same formula for $\widehat{\boldsymbol{Y}}$ applies, replacing $\mathbf{C}$ and $\mathbf{Y}_{a}$ with the extended versions:
[…]
    where
    $\widetilde{\mathbf{C}}=\begin{bmatrix}\mathbf{C}\\ \mathbf{H}\end{bmatrix},\quad\widetilde{\mathbf{Y}}_{a}=\begin{bmatrix}\mathbf{Y}_{a}\\ \mathbf{d}\end{bmatrix}$.
    This augmentation reduces uncertainty in units with weak auxiliary signal and can substantially improve local accuracy.
The estimator $\widehat{\boldsymbol{Y}}$ corresponds to the best linear unbiased predictor (BLUP) under the aggregation constraint. Specifically, it arises from the conditional expectation $\mathbb{E}(\mathbf{Y}|\mathbf{Y}_{a})$ when the latent disaggregated vector $\boldsymbol{Y}$ follows $\mathbb{E}(\boldsymbol{Y})=\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}$ and $\operatorname{Cov}(\boldsymbol{Y})=\mathbf{B}$, and the observed aggregate satisfies $\mathbf{Y}_{a}=\mathbf{C}\boldsymbol{Y}$. Standard results on multivariate normal conditional distributions imply that (Harvey, 1990; Durbin and Koopman, 2012):
$\mathbb{E}(\mathbf{Y}|\mathbf{Y}_{a})=\tilde{\boldsymbol{Y}}+\operatorname{Cov}(\boldsymbol{Y},\mathbf{Y}_{a})\operatorname{Var}(\mathbf{Y}_{a})^{-1}(\mathbf{Y}_{a}-\mathbb{E}(\mathbf{Y}_{a})),\quad\text{with}\quad\tilde{\boldsymbol{Y}}=\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}.$
Substituting the model components yields the explicit form:
$\widehat{\boldsymbol{Y}}=\widehat{\mathbf{A}}^{-1}\mathbf{Z}\hat{\boldsymbol{\beta}}+\widehat{\mathbf{B}}\,\mathbf{C}^{\top}\big{(}\mathbf{C}\widehat{\mathbf{B}}\,\mathbf{C}^{\top}\big{)}^{-1}\Big{(}\mathbf{Y}_{a}-\mathbf{C}\widehat{\mathbf{A}}^{-1}\mathbf{Z}\hat{\boldsymbol{\beta}}\Big{)}$.
By construction, this predictor ensures aggregation coherence since:
Therefore, the predicted disaggregated series exactly sums to the observed total. This guarantees coherence between the disaggregated estimates and the aggregate data while minimizing the conditional prediction variance (Harvey, 1989; Durbin and Koopman, 2012).
Finally, basic regularity conditions are assumed to ensure consistency and asymptotic normality of the estimators, with additional conditions detailed in the following section.
@@
The $\{\epsilon_{it}\}$, $i=1,\ldots,n$ and $t=1,\ldots,T$ are i.i.d with mean zero and variance $\sigma^{2}$. Its moment $\mathbb{E}(|\epsilon_{it}|^{4+\gamma})$ for some $\gamma>0$ exits.
@@
The elements $w_{ij}$ satisfies that $\sum_{j=1}^{n}w_{ij}=1$, $i=1,\ldots,n$. And $\exists j\neq j^{\prime}$ such that:
@@
$n>2$ is fixed for all $t=1,\ldots,T$, $|\rho|<1$ and, $|\phi_{1}|<1$

### v2

@@ Abstract
Spatial models , Spatial autoregressive model , Heteroscedasticity , Spatial regression , Maximum likelihood estimation
@@ 1 Introduction
Spatiotemporal data modeling has gained increasing relevance in econometrics and regional science due to the simultaneous presence of spatial spillovers and temporal dynamics in key macroeconomic indicators. Spatial autoregressive (SAR) models have long been used to capture spatial autocorrelation between neighboring regions, improving inference and prediction by explicitly considering spatial dependencies LeSage and Pace (2015); Cressie and Wikle (2011). Recent developments have extended these models to incorporate temporal structure, leading to comprehensive spatiotemporal frameworks that model both forms of dependence jointly Gelfand et al. (2019); Hsu et al. (2021).
@@
This article contributes to the literature by proposing a new methodology for the spatiotemporal disaggregation of economic aggregates, combining spatial autoregressive models, benchmarking restrictions, and covariate information. Before presenting our approach, it is important to highlight several features that make it attractive both theoretically and practically. First, the method is based on a rigorous asymptotic theory: we establish global identifiability and asymptotic normality of the joint estimator. Second, we introduce a novel anchoring mechanism that incorporates partially observed regional values as linear restrictions. This step mimics real-life situations with sporadic regional data (e.g., for capital cities or large communities) and helps reduce the variance of the predictions. Third, we conducted an exhaustive simulation study encompassing 74,088 parameter configurations, covering a wide range of spatial autocorrelations, covariate informativeness, and grid sizes. These simulations confirmed the robustness and stability of our approach across various scenarios. Finally, we demonstrated the practical relevance of our method with a case study on the disaggregation of Spanish GDP into the 17 autonomous communities between 2002 and 2023, using a large set of auxiliary indicators and principal component analysis for dimensionality reduction.
@@ 2.1 Definition of the model
Let $Y_{t}$ be a time series measured at time $t=1,\ldots,T$ over the spatial unit $\mathcal{R}$, which is subdivided in $n$ spatial regions $\{\mathcal{R}_{1},\ldots,\mathcal{R}_{n}\}$; let $\mathbf{z}_{it}$ a set of $k$ known explicative variables measured at time $t$ in spatial region $i$. Let $Y_{it}$ be the $i$th unknown time serie disagregated at time $t$ over the region $i$ with the restriction defined as:
$\sum_{i=1}^{n}Y_{it}=Y_{t},\qquad\forall t=1,\ldots,T.$ (1)
@@
The objective is to obtain estimated values for $Y_{it}$ using the information allocated both in $\mathbf{z}_{it}$ and $\mathbf{Y}_{a}=(Y_{1},\ldots,Y_{T})$. Let $\boldsymbol{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{k})^{\top}$ be the associate vector of parameters effects of $\mathbf{z}_{it}$ over $\mathbf{Y}$, a spatial autoregressive model is supposed for the vector $\mathbf{Y}_{t}=(Y_{1t},\ldots,Y_{nt})$, the completed model is given by:
$Y_{it}=\rho\sum_{j=1}^{n}w_{ij}Y_{jt}+\mathbf{z}_{it}^{\top}\boldsymbol{\beta}+u_{it}.$ (2)
Let
$\mathbf{Z}$ $=\begin{bmatrix}\mathbf{z}_{11}^{\top}\\ \mathbf{z}_{21}^{\top}\\ \vdots\\ \mathbf{z}_{n1}^{\top}\\ \mathbf{z}_{12}^{\top}\\ \vdots\\ \mathbf{z}_{n2}^{\top}\\ \vdots\\ \mathbf{z}_{nT}^{\top}\\ \end{bmatrix}_{nT\times k},\;\mathbf{Y}=\begin{bmatrix}Y_{11}\\ Y_{21}\\ \vdots\\ Y_{n1}\\ Y_{12}\\ \vdots\\ Y_{n2}\\ \vdots\\ Y_{nT}\\ \end{bmatrix}_{nT\times 1},\;\mathbf{W}=\begin{bmatrix}w_{11}&\cdots&w_{1n}\\ \vdots&\ddots&\vdots\\ w_{n1}&\cdots&w_{nn}\\ \end{bmatrix},\;\mathbf{U}=\begin{bmatrix}u_{11}\\ u_{21}\\ \vdots\\ u_{n1}\\ u_{12}\\ \vdots\\ u_{n2}\\ \vdots\\ u_{nT}\\ \end{bmatrix}_{nT\times 1}.$ (3)
and $\mathbf{A}=\mathbb{I}_{nT}-\rho(\mathbb{I}_{T}\otimes\mathbf{W})=\mathbb{I}_{T}\otimes(\mathbb{I}_{n}-\rho\mathbf{W})$ and $\mathbb{I}_{T}$ the $T$-dimensional identity matrix. Then, the model defined in Equation (2) can be expressed as:
$\mathbf{Y}$ $=\rho(\mathbb{I}_{T}\otimes\mathbf{W})\mathbf{Y}+\mathbf{Z}\boldsymbol{\beta}+\mathbf{U}$
$\mathbf{Y}-\rho(\mathbb{I}_{T}\otimes\mathbf{W})\mathbf{Y}$ $=\mathbf{Z}\boldsymbol{\beta}+\mathbf{U}$
$(\mathbb{I}_{nT}-\rho(\mathbb{I}_{T}\otimes\mathbf{W}))\mathbf{Y}$ $=\mathbf{Z}\boldsymbol{\beta}+\mathbf{U}$
$\mathbf{Y}$ $=\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}+\mathbf{A}^{-1}\mathbf{U}.$ (4)
@@
Let $\mathbf{U}_{i}=(u_{i1},\ldots,u_{iT})$ be an $ARIMA(1,0,0)$, i.e. $\phi_{1}(B)u_{it}=\epsilon_{it}$ This proposal models the unobservable time series analogously to the proposal of Proietti (2006), Quilis (2018), and Bisio (2024). This assumption over $\mathbf{U}_{i}$ allows us to define the covariance matrix of $\mathbf{U}_{i}$ like Sherman (2023), that is:
$\boldsymbol{\Sigma}_{U}=Var(\mathbf{U}_{i})=\frac{\sigma^{2}}{1-\phi_{1}^{2}}\begin{pmatrix}1&\phi_{1}&\phi_{1}^{2}&\cdots&\phi_{1}^{T-1}\\ \phi_{1}&1&\phi_{1}&\cdots&\phi_{1}^{T-1}\\ \phi_{1}^{2}&\phi_{1}&1&\cdots&\phi_{1}^{T-1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \phi_{1}^{T-1}&\phi_{1}^{T-2}&\phi_{1}^{T-3}&\cdots&1\\ \end{pmatrix}$. (5)
@@
where $\phi_{1}$ is the autoregresive parameter of the $ARIMA(1,0,0)$ and $\sigma^{2}$ their associated variance. The problem is that $\mathbf{Y}$ and $\mathbf{U}$ are unobservable by the characteristics of the measurements. But $\mathbf{Z}$ and $\mathbf{Y}_{a}$ are observable and known, then, we define the matrix $\mathbf{C}$ as:
$\mathbf{C}=\mathbb{I}_{T}\otimes\mathbf{1}_{n}^{\top},$ (6)
where $\mathbf{1}_{n}^{\top}$ is a row matrix of ones, such that
$\mathbf{CY}$ $=\mathbf{Y}_{a}=\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}+\mathbf{C}\mathbf{A}^{-1}\mathbf{U}.$ (7)
We obtain that:
$\mbox{Var}(\mathbf{U})_{nT\times nT}$ $=\boldsymbol{\Sigma}_{U}\otimes\mathbb{I}_{n},$
$\mbox{Var}(\mathbf{C}\mathbf{A}^{-1}\mathbf{U})_{T\times T}$ $=\mathbf{C}\mathbf{A}^{-1}(\boldsymbol{\Sigma}_{U}\otimes\mathbb{I}_{n})(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}=\mathbf{C}\mathbf{B}\mathbf{C}^{\top}=\boldsymbol{\Sigma}_{\mathbf{Y}_{a}},$ (8)
$\mbox{Cov}(\mathbf{Y},\mathbf{Y}_{a})_{nT\times T}$ $=\mbox{Cov}(\mathbf{Y},\mathbf{CY})=\mbox{Var}(\mathbf{Y},\mathbf{Y})\mathbf{C}^{\top}=\mathbf{B}\mathbf{C}^{\top},$ (9)
where $\mathbf{B}=\mathbf{A}^{-1}(\boldsymbol{\Sigma}_{U}\otimes\mathbb{I}_{n})(\mathbf{A}^{-1})^{\top}=\boldsymbol{\Sigma}_{\mathbf{Y}}$.
2.2 Estimation and Inference
To develop the maximum likelihood theory, we will first make assumptions of normality and homoscedasticity regarding the model innovations.
@@ 2.2 Estimation and Inference
The innovations satisfy $\epsilon_{it}\sim\mathcal{N}(0,\sigma^{2})$ for all $i=1,\ldots,n$ and $t=1,\ldots,T$.
Under Assumption 1, properties of normal distribution (Harvey (1990), Durbin and Koopman (2012)), and Equations (8) and (9) it follows that
$\mathbf{U}$ $\sim\mathcal{N}_{nT}(\mathbf{0},\boldsymbol{\Sigma}_{U}\otimes\mathbb{I}_{n}),$
$\mathbf{Y}$ $\sim\mathcal{N}_{nT}(\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta},\mathbf{B}),$
$\mathbf{Y}_{a}$ $\sim\mathcal{N}_{T}(\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta},\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}),$ (10)
$\begin{pmatrix}\mathbf{Y}\\ \mathbf{Y}_{a}\end{pmatrix}$ $\sim\mathcal{N}_{nT+T}\left(\begin{pmatrix}\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}\\ \mathbf{C}\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}\end{pmatrix},\begin{pmatrix}\mathbf{B}&\mathbf{BC}^{\top},\\ \mathbf{C}\mathbf{B}&\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}\end{pmatrix}\right)$, (11)
$\mathbf{Y}\mid\mathbf{Y}_{a}$ $\sim\mathcal{N}_{nT}(\boldsymbol{\mu}+\mathbf{BC}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a}),\mathbf{B}-\mathbf{BC}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{CB}),$ (12)
where $\boldsymbol{\mu}=\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}$, and $\boldsymbol{\mu}_{a}=\mathbf{C}\boldsymbol{\mu}$. Therefore, by Equation (10), the generalized least squares estimator for $\boldsymbol{\beta}$ is given by:
$\hat{\boldsymbol{\beta}}$ $=(\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{C}\mathbf{A}^{-1}\mathbf{Z})^{-1}\mathbf{Z}^{\top}(\mathbf{A}^{-1})^{\top}\mathbf{C}^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}\mathbf{Y}_{a}.$ (13)
It should be noted that the estimator $\hat{\boldsymbol{\beta}}$ depends on the values of the parameters $\rho$, $\sigma^{2}$, and $\phi_{1}$. To estimate these parameters, from Equation (10), we have that the density function of $\mathbf{Y}_{a}$ is given by:
$f_{\mathbf{Y}_{a}}(\mathbf{y})=(2\pi)^{-\frac{T}{2}}|\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}|^{-\frac{1}{2}}\exp\left(-\frac{1}{2}(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})\right),\text{ for }\mathbf{y}\in\mathbb{R}^{T},$
then, we solve the maximization problem over the log-likelihood function
$(\hat{\rho},\hat{\sigma}^{2},\hat{\phi}_{1})=\underset{(\rho,\phi_{1},\sigma^{2})\in\mathcal{U}}{\mbox{argmax}}\Bigg\{-\frac{1}{2}\log|\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}|-\frac{1}{2}(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})^{\top}\boldsymbol{\Sigma}_{\mathbf{Y}_{a}}^{-1}(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})\Bigg\}$ (14)
where $\mathcal{U}=(-1,1)\times(-1,1)\times(0,\infty)$.
In practice, solving the problem of Equation (14) can be done using optim function from software R Core Team (2024) with the ”L-BFGS-B” algorithm, which allows the constraints on the parameter space $\mathcal{U}$ to be explicitly imposed. The L-BFGS-B algorithm is a limited-memory version of the Broyden–Fletcher–Goldfarb–Shanno (BFGS) quasi-Newton method, adapted for large-scale problems with simple box constraints Byrd et al. (1995).
To ensure that the parameters are estimable and estimators are sufficient, consistent, and asymptotically normal, the following conditions are proposed:
@@
The elements $w_{ij}$ satisfy that $\sum_{j=1}^{n}w_{ij}=1$ for all $i=1,\ldots,n$. And exist at least one $j\neq j^{\prime}$ such that:
@@
Assumption 2 ensures that at least one spatial unit has a difference between what its neighbors contribute to it and what it contributes to its neighbors in spatial information. This prevents the degenerate case where regions are indistinguishable from one another, yielding multiple solutions. Additionally, the matrix is standardized by rows to ensure that the parameter $\rho$ falls between -1 and 1, facilitating a more accurate interpretation of the autoregressive parameter. This condition is discussed usually in SAR models (Lee, 2004; Kelejian and Prucha, 2010)
@@
The number of spatial regions $n>2$ is fixed for all $t=1,\ldots,T$, $|\rho|<1$ and, $|\phi_{1}|<1$.
Assumption 3 ensures a minimum number of spatial regions and bounds $\rho$ and $\phi_{1}$ to guarantee the stationarity of the ARIMA process.
Assumption 4.
The matrix $(\mathbf{C}\mathbf{A}^{-1}\mathbf{Z}:\mathbf{C}\mathbf{A}^{-1}(\mathbb{I}_{T}\otimes\mathbf{W})\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta})$ has full column rank, that is, the $k+1$ columns are linearly independent.
Assumption 4 is a natural translation of Assumption 8 of Lee (2004) in the context of our model. This assumption ensures that both $\boldsymbol{\beta}$ and $\rho$ are identifiable. Notice that $T>k+1$ is assumed.
This condition is generally satisfied when the number of time periods $T$ greatly exceeds the number of parameters $k+1$, i.e., $T>>k+1$. Intuitively, each time block provides an independent observation of the effect of the spatial operator embedded in $\mathbf{A}^{-1}$ acting on the regressors $\mathbf{Z}_{t}$.